40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 6 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  16819   Thu Apr 28 18:07:04 2022 AnchalUpdateBHDRestoring arm alignment at EOD

Restored arm algiment to get 0.8 max flashing on YARM and 1 max flashing on XARM. I had to move input alignment using TT2-PR3 pair and realign YARM cavity axis using ITMY-ETMY pair.

I would like to advertise this useful tool that I've been using for moving cavity axis or input beam direction. It's a simple code that makes your terminal kind of videogame area. It moves two optics together (either in same direction or opposite direction) by arrow key strokes (left, right for YAW, up, down for PIT). Since it moves two optics together, you actually control the cavity axis or input beam angle or position depening on the mode.

 

  16818   Thu Apr 28 17:45:53 2022 AnchalUpdateBHDPOX Beam issues

[Anchal, Paco]

We investigated the low power issue with POX11 photodiode.

  • We used a power meter to confirm that above 12 uW of power is reaching the ITMX oplev table.
  • But the power reaching the photodiode was only 3 uW, because the 2 steering mirrors used were discarding half of the light. These mirrors were taken from former POP setup and are probably BS or meant for different incidence angle/polarization.
  • We used flashlight to test that the photodiode gives a response, so it is not dead.
  • We also put in a db15 breakout board in line to the PD and confirmed that power input is correct, temperature is nominal, enable is ON, and DC out is responsive to flashlight.
  • So we decided to redo the path with Y1-45P 1inch mirrors. I found two such mirrors in the lab.
  • When setting up the path again, I realized that the beam shape is not even. I took a photo of the beam on the card (see attachment 1) and indeed the POX beam that is coming to the table is half clipped.
  • So tomorrow, we'll open the chamber and find out where this is happening. For now, I've setup a nominal path to the POX11 photodiode, but we'll tune it after we get a clear POX beam on the table.
Attachment 1: PXL_20220429_002703686.jpg
PXL_20220429_002703686.jpg
  16817   Thu Apr 28 11:53:10 2022 AnchalSummaryBHDPOP_SM4 and POP_SM5 Assembly

POP_SM4

I tried out this stack today and found some change of plans.

  • The attachment in elog 40m/16640 says to use 1/4-20 silver coated non-vented screws for joining BA2V to PLS-T238, however PLS-T238's bottom hole is a blind hole and is not vented. So we actually required <1 in long silver plated vented 1/4-20 cap screw for this purpose. Jordan and I were only able to find the correct length silver plated screw but it is not vented. So we decided to make a venting hole on the post from the side.
  • I had to use a 0.14" spacer as washer between PLS-T238 and BA1V. The 1" post shim that Koji got for this purpose had too small a hole in the center to let the 8-32 screw pass (I know, weird). but I think we had 3 spare 0.14" ad we bought 10 when we required 7, so we should be good.
  • The attachment in elog 40m/16640 also says to use 8-32 silver coated vented setscrew for joining TR-1.5 to LMR1V mount. I found one vented silver coated set screw nearby in the clean room but it turned out to be too long. Worse, I overtightened the setscrew when trying this connection which damaged the inner threads of one of the LMR1V. So we need to buy one more LMR1V (maybe an additional spare too) for future installation of OMC1R1/OMC2R1.
    • Then Jordan and I searched for a smaller silver coated and vented 8-32 setscrew but didn't find any. Jordan also noted that LMR1V is an aluminium mount and we should not use silver coated setscrew with it. Since the TR-1.5 mount is ok to be sacrificed if a cold weld happens, we'll just use an uncoated SS 8-32 setscrew to join TR-1.5 and LMR1V. We could not such vented setscrews, but we have plenty of non-vented once. So Jordan is going to make a venthole in TR1.5 top end as well. LMR1V already has a vent hole on its side.

tl,dr; Jordan is preparing PLS-T238 and TR-1.5 with venting holes and C&B and they would be ready by tomorrow. I have collected all other parts for assembly, still looking for the mirror but I know other lab members know where it is, so no big issue there.


POP_SM5

The assemly of this mirror is complete. A slight change here as well, we were supposed to use the former POYM1 (Y1-2037-0) mirror for POP_SM5 but I could not find it. It was stored on the right most edge of the table (see 40m/16450), but it is not there anymore. I found another undocumented mirror on the flow bench on the left edge marked (2010 July: Y1-LW1-2037-UV-0-AR) which means this mirror has a wedge of 1 degree and an AR coating as well. We do not need or care about the wedge or AR coating, so we can use this mirror for POP_SM5. Please let me know if someone was saving this mirror for some other purpose.


I'll finish assembly of POP_SM4 tomorrow and install them in ITMX chamber and resurrect POP path.

Quote:

Here is more detail of the POP_SM4 mount assembly.

It's a combination of BA2V + PLS-T238 + BA1V + TR-1.5 + LMR1V + Mirror: CM254-750-E03
Between BA1V and PLS-T238, we have to do a washer action to fix the post (8-32) with a 1/4-20 slot. Maybe we can use a 1" post shim from thorlabs/newport.
Otherwise, we should be able to fasten the other joints with silver-plated screws we already have/ordered.

I think TR-1.5 (and a shim) has not been given to Jordan for C&B. I'll take a look at these.

 

  16816   Thu Apr 28 09:12:18 2022 AnchalUpdateBHDRestoring arm alignment

This would be a daily first task in the morning. We'll need to check the status of arm alignment and optimize it back to maximum every morning for the rest of the day's work.

Today, when I came, on openin gthe PSL shutter, IMC was aligned good, both arms were flashing but YARM maximum transmission count was around 0.7 (as opposed to 1 from yesterday) and XARM maximum transmission count was 0.5 (as opposed to 1 from yesterday). I did not change the input alignment to the interferometer. I only used ITMY-ETMY to regain flashing count of 1 on YARM and used BS and tehn ITMX-ETMX to regain flashing count of 0.9 to 1 in XARM.

Even thought the oplevs were centered yesterday, I found the oplev had drifted from the center and the optimal position also is different for all ooptics except EMTY and BS. It is worth nothign that in optimal position both PIT and YAW of ITMY and ITMX are off by 70-90 uradians and ETMX Pit oplev is off by 55 uradians.

 

  16815   Wed Apr 27 16:28:57 2022 AnchalSummaryBHDBHD Upgrade - Retreiving arm cavity alignment

[Anchal, Paco, JC]

We had to open ITMY, ETMY chamber doors to get the cavity aligned again. Once we did that, we regained cavity flashing and were able to align the input injection and cavity alignment to get transmission flashing to 1.0 (C1:LSC-TRY_OUT_DQ). JC later centered both ITMY and ETMY oplevs. The ITMY oplev had become completely out of range.

We also opened ITMX, ETMX chamber doors to get Xarm alignment. Again, it seems that ITMX had moved a lot due to cable post installation.

To be continued

  16814   Wed Apr 27 10:05:55 2022 JCUpdateCoil DriversCoil Drivers Update

18 (9 pairs) Coil Drivers have been modified. Namely ETMX/ITMX/ITMY/BS/PRM/SRM/MC1/MC2/MC3.

 

ETMX Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        S2100624                     ETMX Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100631

ITMX Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100620                      IMTX Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100633

ITMY Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100623                   ITMY Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100632

BS Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3            S2100625                     BS Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100649

PRM Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100627                      PRM Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100650

SRM Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100626                        SRM Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100648

MC1 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3         S2100628                        MC1 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3      S2100651

MC2 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3       S2100629                        MC2 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3         S2100652

MC3 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        S2100630                        MC3 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3         S2100653

 

 

Will be updating this linking each coil driver to the DCC

  16813   Tue Apr 26 16:23:22 2022 TommyUpdateElectronicsRFSoC2x2 MTS

We connected a 8 MHz signal generator to the device in order to sync up the ADCs and DACs and hopefully get phase data. 

Some things to note:

  • RF Manual (143)- Need to use XRFDC SYSREF for update event
  • RF Manual (171)- Synchronization steps require us to first enable all clocks and sysref generators (via xrfdc package)
  • RF Manual (173)- Sysref requirments, not clear if PL is syncing as needed.
  • RF Manual (181)- XRFDC example code, see also https://github.com/Xilinx/embeddedsw/blob/master/XilinxProcessorIPLib/drivers/rfdc/examples/xrfdc_mts_example.c

Xilinx RF Manual: https://docs.xilinx.com/v/u/2.4-English/pg269-rf-data-converter

  16812   Mon Apr 25 18:00:03 2022 Ian MacMillanUpdateUpgradeCable supports update

I have designed new cable supports for the new ribbon cables running up the side of the tables in the vacuum chambers. 

The clamps that I have designed (shown in basic sketch attachment 1) will secure the cable at each of the currently used cable supports. 

The support consists of a backplate and a frontplate. The backplate is secured to the leg of the table using a threaded screw. The frontplate clamps the cable to the backplate using two screws: one on either side. Between two fascinating points, the cable should have some slack. This should keep the cable from being stiff and help reduce the transfer of seismic noise to the table. 

It is possible to stack multiple cables in one of these fasteners. Either you can put two cables together and clamp them down with one faceplate or you can stack multiple faceplates with one cable between each faceplate. in this case the stack would go backplate then cable then faceplate then cable then the second faceplate. this configuration would require longer screws.

The exact specifics about which size screws and which size plates to use still have not been measured by me. But it will happen

Attachment 1: Chamber_Leg_Ribbon_Cable_Attachments.pdf
Chamber_Leg_Ribbon_Cable_Attachments.pdf
  16811   Mon Apr 25 17:24:06 2022 AnchalUpdateCDSDAQ still down

I investigated this issue today. At first, it seemed that only new suspension testpoints are inaccessible. I was able to use diaggui for a measurement on MC2. The DAQ network cable between 1X4 and 1Y1 was tied and is very taught now (we should relieve this as soon as possible, best solution is to lay down a longer cable over the bridge). My hypothesis is that the DAQ network might have broken while tying this cable and it probably did not come back since then.

The simplest solution would have been to restart c1su2 models. As I restarted those models though, I found that c1lsc and c1sus models failed. This is very unusual as c1su2 models are independent and share nothing with the other vertex models. So I had to restart all the FE computers eventually. But this did not solve the issue. Worse, now the DAQ isn't working for the vertex machiens as well.

Next step was to try restarting fb1 itself. We switched off all the FE computers, restarted fb1, stopped daqd_* processes, reloaded gpstime module, restarted open-mx, mx, nds and daqd_* process. But the mx.service failed to load with following error message:

● mx.service - LSB: starts MX driver for Myrinet card
   Loaded: loaded (/etc/init.d/mx)
   Active: failed (Result: exit-code) since Mon 2022-04-25 17:18:02 PDT; 1s ago
  Process: 4261 ExecStart=/etc/init.d/mx start (code=exited, status=1/FAILURE)

Apr 25 17:18:02 fb1 mx[4261]: Loading mx driver
Apr 25 17:18:02 fb1 mx[4261]: insmod: ERROR: could not insert module /opt/mx/sbin/mx_mcp.ko: File exists
Apr 25 17:18:02 fb1 mx[4261]: insmod: ERROR: could not insert module /opt/mx/sbin/mx_driver.ko: File exists
Apr 25 17:18:02 fb1 systemd[1]: mx.service: control process exited, code=exited status=1
Apr 25 17:18:02 fb1 systemd[1]: Failed to start LSB: starts MX driver for Myrinet card.
Apr 25 17:18:02 fb1 systemd[1]: Unit mx.service entered failed state.

(Ignore the timestamp above, I ran the command again to capture the error message.)

However, I was able to start all the FE models without any errors and daqd processes are also all running without showing any errors. Everything is green in CDS screen with no error messages. But the only thing still wrong is mx.service which is not running.

From my limited knowledge and experience, mx.service is a one-time script that mounts mx devices in /dev and loads the mx driver. I tried running the script /opt/mx/sbin/mx_start_stop :

controls@fb1:/opt/mx/sbin 1$ sudo ./mx_start_stop start
Loading mx driver
insmod: ERROR: could not insert module /opt/mx/sbin/mx_mcp.ko: File exists
insmod: ERROR: could not insert module /opt/mx/sbin/mx_driver.ko: File exists

This gave the same error. On searching little bit online, "insmod: ERROR; cound not insert module" error comes up when the kernel version of the driver doesnot match the Linux kernel (whatever that means!). Such deep issues should not appear out of nowhere in a previosuly perfectly runnig system. I'll check around more what changed in fb1, network cables etc.

  16810   Mon Apr 25 16:57:57 2022 AnchalUpdateBHDCable post installation

[Anchal, Tega, JC]

We installed cable posts in ITMY, BS, and ITMX chambers for all the new suspensions. Now, there is no point where the OSEM connections are hanging freely.

In BS chamber, we installed one post for LO2 near the north edge of the table and another post for PR3 on the Western edge with the blue cable running around the table on the floor.

In ITMY chamber, we installed the cable post in between AS1 and AS4 with the blue cables running around the table on the floor. This is to ensure the useful part of the table remains empty for future and none of the OSEM cables are taught in air.

 

  16809   Mon Apr 25 14:49:02 2022 KojiUpdateGeneralNitrogen Tank

For your (and mine) info:

N2 pressure can be monitored on the 40m summary page: https://nodus.ligo.caltech.edu:30889/detcharsummary/day/20220425/vacuum/
(you need to hit "today" to go to the current status)

 

  16808   Mon Apr 25 14:19:51 2022 JCUpdateGeneralNitrogen Tank

Coming in this morning, I checked on the Nitrogen tanks to check the level. One of the tanks were empty, so I went ahead and swapped it out. One tank is at 946 PSI, the other is at 2573 PSI. I checked for leaks and found none.

  16807   Sun Apr 24 13:17:08 2022 TommyUpdateElectronicsNew RFSoC2x2 Overlay

We recieved an overlay from Chris Stoughton which he used for a ZCU11 board. The overlay is supposed to be compatible with the RFSoC 2x2 and help enable the Multi-Tile Synchronization (MTS) we need. He also provides a .py with the necessary low level connection to the board and its memory along with a few sample notebooks.

Progress So Far:

  • The overlay loads properly and in reasonable time. 
  • We can set the mixer and dac frequencies. However, it is unclear what event_src Chris wanted for the adc mixers. It seems that he was using event_src_immediate (possibly unintenionally) which is not an available adc mixer setting in our board. Instead, we set the event source to "Tile" and will later determine if this is an issue.
  • We then go to get data from the buffer. There are two functions called: capture and transfer. These are called on the pynq DMA. Capture runs fine but we get stuck during the transfer at dma.revchannel.wait(). This issue has not been resolved.
  16806   Fri Apr 22 14:14:33 2022 JordanUpdateVACTP3 Forepump tip seal replacement

Jordan, JC

While the pumpspool is vented, I thought it would be a convenient time to change out the tip seal on the TP3 forepump. This one had not been changed since 2018, so as preventative maintence I had JC remove the pump and begin cleaning/installing the new tip seal.

Unfortunately the tip seal broke, but I have ordered another. We should have this pump ready to go late next week. If one is needed sooner, there is a spare IDP 7 pump we can install as the TP3 forepump.

Attachment 1: IMG_0572.jpeg
IMG_0572.jpeg
Attachment 2: IMG_0573.jpeg
IMG_0573.jpeg
  16805   Fri Apr 22 12:15:08 2022 AnchalUpdateBHDCable post installation

If someone gets time, let's put in all the cable posts and clean up our cable routing on the tables.

 

  16804   Fri Apr 22 12:09:51 2022 AnchalUpdateCoil DriversPlease update DCC pages

Nice. Please put this information on the DCC pages of the coil driver units. You'll find links to all the units in this document tree LIGO-E2100447. For each page, click on "Change Metadata" from the left panel and add the change made to the resistor (including the resistor name on PCB, previous and new value), and add a link to your previous elog post which has more details like photos, to "Notes and Changes", and upload an updated version of the circuit schematic by creating an annotation in the previous circuit schematic pdf. Every unit that has a serial number in the lab has a DCC page (if not, we should create one) where we should track all such hard changes.

 

  16803   Fri Apr 22 12:03:09 2022 TegaUpdateBHDAS1 UR OSEM problem localized in the chamber

[JC, Tega, Ian, Paco]

We found that the UR cable was clamped to the table by one of the ITMY OPLEV steering mirror mounts that was recently installed. After freeing the cable, the UR signal is now active again.

Quote:

This indicates that the AS1 UR OSEM problem is localized in the chamber. Please check if the DSUB pins are touching the table or something else.

 

  16802   Fri Apr 22 07:01:58 2022 JcUpdateCoil DriversAdding Resistors and Reinstalling

[Koji, JC]

Coil Drivers LO2, SR2, AS4, and AS1 have been updated a reinstalled into the system. 

LO2 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        Unit: S2100008

LO2 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3                    Unit: S2100530

SR2 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        Unit: S2100614

SR2 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3                    Unit: S2100615

AS1 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        Unit: S2100610

AS1 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3                    Unit: S2100611

AS4 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        Unit: S2100612

AS4 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3                    Unit: S2100613

Attachment 1: IMG_0536.jpeg
IMG_0536.jpeg
Attachment 2: IMG_0539.jpeg
IMG_0539.jpeg
Attachment 3: IMG_0542.jpeg
IMG_0542.jpeg
Attachment 4: IMG_0545.jpeg
IMG_0545.jpeg
Attachment 5: IMG_0548.jpeg
IMG_0548.jpeg
Attachment 6: IMG_0551.jpeg
IMG_0551.jpeg
Attachment 7: IMG_0554.jpeg
IMG_0554.jpeg
Attachment 8: IMG_0557.jpeg
IMG_0557.jpeg
  16801   Thu Apr 21 20:33:31 2022 KojiUpdateBHDAS1 UR OSEM problem localized in the chamber

Tega and Paco reported that the UR OSEM of AS1 lost the response.

- I have checked the LED MON (left) of the satellite amp for AS1. CH1/2/3 had 5V -> This indicates that the OSEM LEDs are (most likely) functioning.
- Then I went to the ITMY flange and connected the OSEM emulator instead of the Dsub25 cable. The attachment shows that the UR OSEM LED/PD worked fine with the OSEM emulator. WIth the vacuum flange connected it lost the response.

This indicates that the AS1 UR OSEM problem is localized in the chamber. Please check if the DSUB pins are touching the table or something else.

 

Attachment 1: Screenshot_from_2022-04-21_20-13-13.png
Screenshot_from_2022-04-21_20-13-13.png
  16800   Thu Apr 21 18:42:47 2022 TegaUpdateBHDPart V of BHR upgrade - Align LO and AS beams to BHD BS

[Tega, Yuta, Paco]

We tried aligning the LO and AS beams on to the BHD beamsplitter. During the alignment process, we noticed that the damping loop for AS1 was not working. Paco drew our attention to the fact that the UR OSEM signal was alway close to zero, so we checked to ensure that the magnet was still within the OSEM recess and it looks OK. Next we checked the electrical connection at the interface between the copper OSEM cables to the blue in-vacuum flat cable and this too looks alright also. Since the AS1 coil driver was recently modified, it is possible we might find the problem there, so I'll ask Koji about this.


So Koji clarified that the coil driver board and SATAMP boards are different so we should connect this issue to the coil driver board.

  16799   Thu Apr 21 18:18:42 2022 yehonathanUpdateBHDPOX Alignment

{Yehonathan, Paco}

BS, ITMX and ETMX were aligned to get flashing in the X arm.

I aligned the POX beam on the ITMX table using a mixture of the old POP and POX optics. The beam was stirred to the POX11 RFPD. We measure the DC power using a scope but we see nothing. We went and saw that the POX11 cable was not connected to RF rack so we connected it along with some other RFPD cables.

We return but there is still no DC. We ndscope C1:LSC-POX11_I_ERR_DQ C1:LSC-POX11_Q_ERR_DQ and maximize the signal (attachment). The readout is very weak though. It should be as strong as POY which we already observed to have good SNR.

We also noticed that the one of the beam dumps for the POX RFPD is not glued and easily falls down.

Attachment 1: POX11_alignment.png
POX11_alignment.png
  16798   Thu Apr 21 17:32:35 2022 Ian MacMillanSummarySEISeismic Study of Buildings and Caltech Campus

[Rana, Ian]

We built a power supply for the accelerometer shown in Attachment 1 based on the diagram shown in the Wilcoxon manual and shown in attachment 2. We used a 9V power supply and a capacitor value of 680uF. We did not use a constant current diode. 

When hooked up to an oscilloscope we saw vibrations from hitting our hands on the table but we did not see the same amplitude in the negative and positive directions. For example, when I held the accelerometer and moved it down you would see a dip then a peak as the accelerometer accelerated down then accelerated up when I stopped the down word movement. But weirdly when I did the opposite (moved the accelerometer up the same dip then a peak appeared. This is a little concerning because it should be the opposite. it should be a peak then a dip. This in addition to the seemingly decreased sensitivity in one direction make me think that the accelerometer is broken.

I labeled the box with "might be broken" before I returned it to the cryo lab.

Attachment 1: IMG_1820.jpg
IMG_1820.jpg
Attachment 2: Screen_Shot_2022-04-21_at_5.52.51_PM.png
Screen_Shot_2022-04-21_at_5.52.51_PM.png
  16797   Thu Apr 21 16:49:01 2022 PacoUpdateBHDPart V of BHR upgrade - BS Oplev + LO1 offset

[Paco, JC, Yuta]

We aligned the BS oplev using the new BSOL mirror pair. The main change is now the AOI of the oplev on the BS is quite normal. The output beam in the in-air table was quite large (diverging?) so we had to place a short FL lens in front of the QPD.


Separately, I added the LO1 YAW offset of ~ -2500 counts (before the coil driver changes it was -24500 counts) and saw LO beam hitting LO2. This means the alignment of the LO beam can move downstream.

  16796   Thu Apr 21 16:36:56 2022 TegaUpdateVACcleanup work for vacuum git repo

git repo - https://git.ligo.org/40m/vac

Finally incorporated the FRGs into the main modbusIOC service and everything seems to be working fine. I have also removed the old sensors (CC1,CC2,CC3,CC4,PTP1,IG1) from the serial client list and their corresponding EPICS channels. Furthermore, the interlock service python script has been updated so that all occurrence of old sensors (turns out to be only CC1) were replaced by their corresponding new FRG sensor (FRG1) and a redundnacy was also enacted for P1a where the interlock condition is replicated with P1a being replaced with FRG1 because they both sense the main volume pressure.

  16795   Thu Apr 21 15:22:44 2022 KojiUpdateSUSOutpur resistors updates for SR2, LO2, AS1, and AS4 done

[JC Koji]

Quick report: JC has done all the mods for the coil driver circuit in the morning and we worked on the reinstallation of them in the afternoon.
I'll check the damping loops / sus servo settings. JC is going to make an ELOG entry and DCC updates for more precise record of the mods.

  16794   Thu Apr 21 11:31:35 2022 JCUpdateVACGauges P3/P4

[Jordan, JC]

It was brought to attention during yesterday's meeting that the pressures in the vacuum system were not equivalent althought the valve were open. So this morning, Jordan and I reviewed the pressure gauges P3 and P4. We attempted to recalibrate, but the gauges were unresponsive. Following this, we proceeded to connect new gauges on the outside to test for a calibration. The two gauges successfully calibrated at atmosperic pressure. We then proceeded to remove the old gauges and install the new ones. 

Attachment 1: IMG_0560.jpeg
IMG_0560.jpeg
Attachment 2: IMG_0561.jpeg
IMG_0561.jpeg
  16793   Thu Apr 21 10:35:23 2022 KojiUpdateCDSDAQ seemed down

Yesterday, when I worked on the damping servo, I found that any of the daqvtools (ndscope, dtt, dataviewer,...) is not available.  We may need to restart the fb and rt machines.

  16792   Wed Apr 20 18:50:03 2022 KojiUpdateBHDPart V of BHR upgrade - PR2 weirdness

It seemed that it comes from the servo oscillation. This does not happen when the output limitters were set to be 100-ish. But even so the gains looked quite low.

I turned on the Cheby rool-offs for all the DOFs, and this allowed me to increase the damping gain A LOT.
The gains were 2~5 but now they are now 20-25 for the face OSEMs and 150 for SD.

The attached is the example of the damping when all the damping loops are on.

I think we need to tune the servo loops carefully for all the SUSs by actually looking at the openloop transfer functions rather than a personal feeling. => To Do

Attachment 1: Screenshot_2022-04-20_18-46-11.png
Screenshot_2022-04-20_18-46-11.png
  16791   Wed Apr 20 16:11:08 2022 KojiUpdateSUSOutpur resistors updated for LO1 coil drivers / for SR2, LO2, AS1, and AS4 in progress

[JC Koji]

To give more alignment ranges for the SUS alignment, we started updating the output resistors of the BHD SUS coil drivers.
As Paco has already started working on LO1 alignment, we urgently updated the output Rs for LO1 coil drivers.
LO1 Coil Driver 1 now has R=100 // 1.2k ~ 92Ohm for CH1/2/3, and LO1 Coil Driver 2 has the same mod only for CH3. JC has taken the photos and will upload/update an elog/DCC.

We are still working on the update for the SR2, LO2, AS1, and AS4 coil drivers. They are spread over the workbench right now. Please leave them as they're for a while.
JC is going to continue to work on them tomorrow, and then we'll bring them back to the rack.

Attachment 1: IMG_0527.jpeg
IMG_0527.jpeg
Attachment 2: IMG_0528.jpeg
IMG_0528.jpeg
Attachment 3: IMG_0529.jpeg
IMG_0529.jpeg
Attachment 4: IMG_0530.jpeg
IMG_0530.jpeg
Attachment 5: IMG_0531.jpeg
IMG_0531.jpeg
Attachment 6: IMG_0532.jpeg
IMG_0532.jpeg
  16790   Wed Apr 20 14:56:06 2022 TommyUpdateElectronicsRFSoC 2x2 board -- setup for remote work & BALUN saga

Here are a few options for replacement BALUNs from Mini Circuits and specs:

Current. TCM1-83X+, 10-8000 MHz, 50 Ohms, Impedance Ratio 1, Configuration K

1. Z7550-..., DC-2500 MHz (some DC-2300), 50/75 Ohms, Impedance Ratio 1.5, Configuration Q. There are various types of the Z7550 which have different connectors (SMA and BNCs). These have much larger dimensions than the TCM1-83X. Can handle up to 5A DC current with matching loss 0.6 dB.

2. SFMP-5075+, DC-2500 MHz, 50/75 Ohms, Impedance Ratio 1.5, Configuration D. This is an SMA connected BALUN. It can handle 350mA, has a matching loss 0.4 dB, and has 1W power handling.

Quote:

Seems like it should be possible to just remove the transformer (aka as a BALUN ... BALanced, UNbalanced), or replace it with a lower frequency part. Its just a usual mini-circuits part. Maybe you can ask Chris Stoughton about this and ask Tommy to checkout some of the RFSoC user forums for how to go to DC.

Quote:
 

After fiddling around with the tone-generators and spectrum analyzer tools in loopback configuration (DAC --> ADC direct connection), we noticed that lower frequency (~ 1 MHz) signals were hardly making it out/back into the board... so we looked at some of the schematics found here and saw that both RF data converters (ADC & DAC) interfaces are AC coupled through a BALUN network in the 10 - 8000 MHz band (see Attachment #1). This is in principle not great news if we want to get this board ready for audio-band DSP.

We decided that while Tommy works on measuring TFs for SHP-200 all the way up to ~ 2 GHz (which is possible with the board as is) I will design and put together an analog modulation/demodulation frontend so we can upconvert all our "slow" signals < 1MHz for fast, wideband DSP. and demodulate them back into the audio band. The BALUN network is pictured in Attachment #2 on the board, I'm afraid it's not very simple to bypass without damaging the PCB or causing some other unwanted effect on the high-speed DSP.

 

model_no case_style single2single single2bal bal2bal center_tap dc_iso freq_low freq_high impedance imped_ratio interface tech config
SFMP-5075+ FF1891 Y N N N N DC 2500 50/75 1.5 CON CORE & WIRE D
TCM1-83X+ DB1627 N Y Y N N 10 8000 50 1 SMT CORE & WIRE K
Z7550-BFNF+ H795-14 Y N N N N DC 2500 50/75 1.5 CON CORE & WIRE Q
Z7550-BMBF+ QP1876-1 Y N N N N DC 2300 50/75 1.5 CON CORE & WIRE D1
Z7550-BMNF+ QP1876 Y N N N N DC 2500 50/75 1.5 CON CORE & WIRE Q
Z7550-FFNM+ H795-1 Y N N N N DC 2300 50/75 1.5 CON CORE & WIRE Q
Z7550-FFSF+ H557-1 Y N N N N DC 2500 50/75 1.5 CON CORE & WIRE Q
Z7550-FMSF+ H795-3 Y N N N N DC 2300 50/75 1.5 CON CORE & WIRE Q
Z7550-FMSFDC+ H795-3 Y N N N Y 1 2500 50/75 1.5 CON CORE & WIRE Q
Z7550-NFNF+ H795-10 Y N N N N DC 2500 50/75 1.5 CON CORE & WIRE D1
Z7550-NMNF+ H795-4 Y N N N N DC 2300 50/75 1.5 CON CORE & WIRE Q
  16789   Wed Apr 20 08:20:22 2022 PacoUpdateBHDPart V of BHR upgrade - PR2 weirdness

Yesterday, I tried tuning the PR2 damping gains by increasing the gain until the damping gave the ~5 oscillations (by watching the damped motion using StripTool, and keeping an eye on the PD var). I noticed that often when I changed the gain, some OSEM sensors shifted (gained an offset!!) and the PD var values changed, typically increasing at higher damping gains. I reverted the changes until the PD var looked "normal" again (~ 2 mV) but it is hard to imagine that the damping filters can have such a "DC" effect, given the shape comprises single zero at 0 Hz (and pole at 30 Hz).

  16788   Tue Apr 19 18:10:10 2022 PacoUpdateBHDPart V of BHR upgrade - POX11 path, LO path, and ITMX Oplev

[Yuta, Paco]

We set up POX11 beam path from ITMX chamber to the ITMX in-air table. To do this, we first identified the POX reflection on the ITMX chamber, and then steered the POXM1 (in the BSC) by hand until we cleared the viewport. We also checked that the POX beam is centered on POXM1.

We then decided to slide the LO1 YAW to clear the LO beam path, which was otherwise clipping on the PR2 SOS. The slider (DAC limited) range of -25000 counts was barely enough to clear the SOS comfortably and avoid hitting the POXM1. The LO beam is now hitting LO2 mirror, so LO alignment can proceed from BSC and ITMY chamber.

Finally, we aligned the input ITMX Oplev beam to ITMXOL2, then ITMX, then to ITMXOL2, and finally into the ITMX in-air optical table. We took some photos of the Oplev beam (see Attachments) to note their position.

By the end of in-vacuum work there was still some flashing in the arm cavities, but fine alignment is required.


After closing the ITMX Chamber, and BSC, we moved on to center the ITMXOL beam. We accomplished this by using two mirrors instead of one as was previously the case. This relaxed the angle of incidence a little, but we had to change the path and the position of the QPD. The QPD sum reads ~ 6600 counts versus the ~ >8000 counts it read right before the vent. One attempt at closing the OL loop, and the ITMX starting oscillating in YAW (PIT was ok), so we realized that maybe we flipped the order in which the OL1 / OL2 mirror were arranged and so the YAW loop needed to flip its sign. Indeed after changing the C1:SUS-ITMX_OL_YAW_GAIN from -6.0 to 6.0 the OL_YAW loop is stable.

Attachment 1: DJI_0167.JPG
DJI_0167.JPG
Attachment 2: DJI_0169.JPG
DJI_0169.JPG
Attachment 3: DJI_0168.JPG
DJI_0168.JPG
  16787   Mon Apr 18 23:22:39 2022 KojiUpdateGeneralTool box and Work Station Organization

Whoa! Thanks!

Attachment 1: PXL_20220419_062101907.jpg
PXL_20220419_062101907.jpg
  16786   Mon Apr 18 17:53:45 2022 PacoSummaryBHDPart IIa of BHR upgrade - POY11 debugging

[Paco, Ian]

We aligned the X-arm IR laser

  • First fix pointing on the ITMX by moving the BS (mostly pitch)
  • then open etmx chamber and fine tune the pointing using the BS until it is centered on ETMX
  • using beam card we quickly see the third reflection back misaligned on pitch so we move ITMX pitch to center it on ETMX
  • fine tune the ITMX alignment on ETMX and check higher order reflaction overlaping with input

at this point we checked C1:LSC-TRX_DQ using ndscope and luckly we see a tiny bit of flashing (about 0.04 in normalized high gain PD counts). So we closed ETM chamber and ITM chamber and go to control room to optimize this signal. The optimization was done in the following way:

  • First, just reiterate on the last steps from above, by maximizing the peak transmission using ITMX/ETMX pair.
  • Then, slide BS alignment, mostly in PIT, and return to ITMX/ETMX pair.
    • At some point, turning the BS PIT made a huge improvement, so I turned the control room light off and looked at the camera on the quad monitors.
  • Based on the location of the flashes (now brighter) on the ITMX/ETMX, the beam seemed to be off in PIT more than YAW, so we focused on correcting the pointing (moving BS) and then correcting with ITMX/ETMX, until the flashes got centered around ETMX.

Final peak transmission (C1:LSC-TRX_DQ) was ~ 1.3 in normalized high gain PD counts. Power budgeting tells us the peak should be ~ 2.0. The flashing on this arm is much better than YARM, so we will press on by installing POX11 RFPD and attempt locking tomorrow. This also means that YARM can be improved by a combination of alignment and/or SUS damping.

In the end we turned on the ETMX oplev and centered it to "save" our flashing position using this reference.

  16785   Mon Apr 18 16:09:07 2022 YehonathanUpdateBHDBHD Readout simulation

I'm planning on simulating the BHD readout noise in a manner very similar to the ALS noise model using Simulink. I've made a sketch of the model for the longitudinal DOFs (attached). A model for ASC will be similar but with more measurement devices (OpLevs, QPDs, WFSs).

I'm not pretending to simulate everything in this diagram on the first go, it is just a sketch of the big picture.

Attachment 1: BHD_Simulink_Sketch.drawio.pdf
BHD_Simulink_Sketch.drawio.pdf
  16784   Mon Apr 18 15:17:31 2022 JancarloUpdateGeneralTool box and Work Station Organization

I cleaned up around the 40 m lab. All the Laser Safety Glasses have been picked up and placed on the rack at the entrance.

Some miscellaneous BNC Connector cables have been arranged and organized along the wall parallel to the Y-Tunnel.

Nitrogen tanks have been swapped out. Current tank is at 1200 psi and the other is at 1850 psi.

The tool box has been organized with each tool in its specified area.

  16783   Mon Apr 18 14:52:47 2022 Ian MacMillanSummarySEISeismic Study of Buildings and Caltech Campus

[Ian, JC]

I want to take measurements of seismic noise at different places on Caltech's campus and in different buildings. I will try to use the accelerometer in my phone for this but first I must calibrate it (Against the 40m accelerometers). 

I placed my iPhone 11 pro next to the seismometers at the 40m MC as seen in Attachment 1.

The calibration from the instrument was done using cts/rthz * 1V/16384cts * 1/ampgain * g/10V * 10m/s^2/g. The ampgain for all was 100.

Next, I took 100 seconds of data on both the iPhone and the three orthogonal Wilcoxon accelerometers.

The ASD for both of the total acceleration is shown in Attachment 2

The ASD for the individual directions acceleration is shown in Attachment 3

The coherence between the individual directions acceleration and the 40m's individual directions is shown in Attachment 4. For this calculation, the 40m data were downsampled to roughly match the phone's sample rate. This coherence is not very good. It should be higher. Because the phone and 40m sensors were picking up the same data as the phone. Because of this I also looked at the coherence between the individual 40m sensors.

In Attachment 5 I look at the coherence between the individual 40m sensors. This should give me a good idea of whether this is some other issue giving me mow coherence. This plot shows that the coherence between the individual 40m sensors is much better than between the phone and the 40m sensors.

Now I wanted to see what kind of data the iPhone could get from real-world tests. I placed it in a number of locations described below and plotted their ASDs in Attachment 6. The locations are thus:

Identifier  Location Notes
QIL QIL Lab in the Sub-basement of west bridge In sub-basement not much activity when taking measurements.
WBSH West bridge sub-basement hallway on floor in hallway no activity around
WB1H West Bridge 1st floor Hall placed on the floor near pillar near stairs to LIGO offices on the ground floor of west bridge
40m desk on my desk at the 40m placed on the desk while people were walking around and I had my feet on the desk. should be noisy

Notice how at the low end the amplitudes follow the relative amplitudes I would expect. QIL and WBSH are the lowest then WB1H is noisier and 40m desk is the noisiest. However, this is only true up until about 0.5 Hz then they all overlap. Since I would expect the 40m desk should be much noisier at all frequencies I suspect that the phone accelerometer is not suitable for measurements higher than 0.5 Hz.

Possible Problems:

One possible problem with my measurement is that my phone was in a leather case. this may have damped out higher frequencies. Also, my phone was not weighed down or bolted to the floor. this stronger connection would make it better at detecting higher frequencies. I could repeat the experiment with no case and a weight on top of my phone.

What's next:

Since I don't think the phone can give me accurate data above 0.5Hz for quiet environments. It may not be suitable for this task. It would seem that the right instrument is the Wilcoxon 731A but it requires an amplifier that I can't track down.

 

I included all the data and code in the zip file in attachment 7

 

Attachment 1: IMG_0513.jpg
IMG_0513.jpg
Attachment 2: tot_acc_cal.pdf
tot_acc_cal.pdf
Attachment 3: indiv_acc_cal.pdf
indiv_acc_cal.pdf
Attachment 4: dec_Coherence.pdf
dec_Coherence.pdf
Attachment 5: 40m_self_Coherence.pdf
40m_self_Coherence.pdf
Attachment 6: tot_acc_testsites.pdf
tot_acc_testsites.pdf
Attachment 7: Calibration.zip
  16782   Fri Apr 15 11:59:16 2022 YehonathanUpdateIOOIMC completely misaligned

Came this morning, opened the PSL and there was not even a beam on the MC REFL.

Looking at the big monitor it seems like the WFS signals went through the roof during the "auto-alignment" night session.

I restored the MC alignment from before the misalignment happen and wait for the SUS to damp. Once the RMS values went below 200 I enabled the watchdog and the coil outputs.

I opened the PSL shutter and the IMC locked immediately. I turned on the WFS servo and the MC REFL DC went down to 0.3. I run the WFS relief script.

  16781   Thu Apr 14 18:39:13 2022 PacoSummaryBHDPart IIa of BHR upgrade - POY11 debugging

[Paco]

In reply to Koji's questions;

Q1: What is the product number of the broken Kepco?

A1: Kepco JQE 0-25 V 4A (1/4 rack mountable 100 W linear power supply)

Q2: Do you feel the burning smell on this broken kepco?

A2: Not very clearly. It just smells like generic broken electronic, but the fan was already long gone by the time we detected its problem.

Q3: How much current does the +15 VDC line draw from the Sorensen?

A3: The Sorensen current monitor reads 6.3 Amps

Q4: Is there a linear power supply in the market that can handle this much current with some margin?

A4: Probably... but we would have to look around.

Q5: Do we really need a linear power supply there?

A5: Good question, I guess anything that is not contaminating the RF electronics with HF noise (e.g. switching PS) can work?

Q6: Is the same fan problem happening on other Kepcos?

A6: Other fans are on, but at least one or two have the "ill" sound... It may be worthwile to give them maintenance if we can.

  16780   Thu Apr 14 18:34:51 2022 PacoSummaryBHDITMY Oplev reinstalled (Re: 2 in oplev mirrors incompatible with LMR2V)

[Paco, Yehonathan]

We installed ITMYOL1 and ITMYOL2 on the ITMY chamber. We aligned the ITMY OpLev beam and closed the loop successfully, we then had a second round of YARM aligment, where we brought the Y peak transmission up from 0.04 counts to 0.09 counts (up by a factor of two). We still couldn't close the YARM loop but we have a better alignment.

  16779   Thu Apr 14 11:52:57 2022 yehonathanSummaryBHDPart IIa of BHR upgrade - POY11 debugging

{JC, Paco, Yehonathan, Ian}

POY lens was moved to infront of the POY steering mirror to make the POU beam focused on the POY11 RFPD. We measured the DC output with an oscilloscope and optimized it with the steering mirrors. We get ~ 16.5mV.

The new lens position blocked the BS OpLev ingoing beam, so we repositioned the OpLev mirrors to make the beam path not hit the lens.

We went to the control room to observe the PDH signal. We observed a series of PDF osscillation and then the signal died infront of our eyes! There is just noise.

We go and check the +/-15V powering the RFPD and we find that the V- is ~ 14V which is good but the V+ was ~ 2.7V which is not.

We went to the PD interface and measure the POY11 output oltages using a breakout board and got the same result.

The PD interface was taken out for inspection. All the OP27 on channel 3 were replaced with new ICs (without need turns out)...


The PD interface card turned out to be OK. What happened is that one of the Kepcos in the RF rack died because its fan crumbled as seen in Attachment #2 (could this be the source of burning smell?). In response, the rack was drawing from the other Kepco (connected in parallel) way too much current  (4A) and the current limiter dropped its voltage from 15V to 2.7V.

The Kepco pair was removed and replaced with a single Sorensen. The POY PDH signal was restored (see attachment).

Attachment 1: Screenshot_2022-04-14_15-53-39.png
Screenshot_2022-04-14_15-53-39.png
Attachment 2: PXL_20220414_220616875.jpg
PXL_20220414_220616875.jpg
  16778   Thu Apr 14 10:18:35 2022 PacoSummaryBHD2 in oplev mirrors incompatible with LMR2V

[Paco, JC]

We realized the 2 in oplev mirrors (Thorlabs BB2-E02) for ITMYOL, SRMOL, and BSOL, are 0.47 in thick, while the LMR2V fixed mount is 0.46 in deep, even without taking the retaining ring into account. After a brief exchange with Koji, and Ian, we decided to glue the mirrors onto the mounts using Torr Seal (a vac compatible epoxy). They are curating in the clean room and should be ready to install in about 2 hours.

  16777   Thu Apr 14 09:04:30 2022 JordanUpdateVACRGA Volume RGA Scans

Prior to venting the RGA volume on Tuesday (4/12/2022) I took an RGA scan of the volume to be vented (RGA+TP1 volume+Manual Gate Valve) to see if there was a difference after replacing the manual gate valve. Attached is the plot from 4/12/22, and an overlay plot to complare 4/12/22 to 12/10/2021, when the same volume was scanned with the old (defective) manual gate valve.

There is a significant drop in the ratio O2 compared the the nitrogen peak and reduced Argon (AMU 40) which indicates there is no longer a large air leak.

12/10/21 N2/O2 ratio ~ 4 (Air 78%N2 / 21%O2)

4/12/22 N2/O2 ratio ~ 10      

There is one significant (above noise level) peak above AMU 46, which is at AMU 58. This could possibly be acetone (AMU 43 and 58) but overall the new RGA Volume scans look significantly better after the manual gate valve replacement. Well done!

Attachment 1: 40mRGA_Overlay.pdf
40mRGA_Overlay.pdf
Attachment 2: RGAVolume_4_12_22.PNG
RGAVolume_4_12_22.PNG
  16776   Wed Apr 13 18:55:54 2022 KojiUpdateCamerasCamera Battery Test

I believe that the Nikon has an exposure problem and that's why we bought the Canon.

 

  16775   Wed Apr 13 16:23:54 2022 Ian MacMillanUpdateGeneralSmell in 40m

[Ian, Paco, JC]

There is a strange smell in the 40m. It smells like a chemically burning smell maybe like a shorted component. I went around with the IR camera to see if anything was unusually hot but I didn't see anything. The smell seems to be concentrated at the vertex and down the y-arm

  16774   Wed Apr 13 15:57:25 2022 Ian MacMillanUpdateCamerasCamera Battery Test

Tested the Nikon batteries for the camera. they are supposed to be 7V batteries but they don't hold a charge. I confirmed this with multi-meter after charging for days. Ordered new ones Nikon EN-EL9

  16773   Wed Apr 13 12:50:07 2022 TegaUpdateVACNew Pressure Gauge Install/Pump Spool Vent

[Jordan, JC, Tega]

We have installed all the FRGs and updated the VAC medm screens to display their sensor readings. The replacement map is CC# -> FRG#, where # in [1..4] and PRP1 -> FRG5. We now need to clean up the C1VAC python code so that it is not overloaded with non-function gauges (CC1,CC2,CC3,CC4,PRP1). Also, need to remove the connection cables for the old replaced gauges.

Attachment 1: VAC_FRG_UPGRADE.png
VAC_FRG_UPGRADE.png
  16772   Tue Apr 12 09:05:21 2022 JordanUpdateVACNew Pressure Gauge Install/Pump Spool Vent

Today, Tega and I would like to vent the pump spool an dinstall the new FRG-400 Agilent Pressure Gauges (per elog 15703). The attached picture shows the volume needed to be vented highlighted in red, and the gauges that need to be replaced/removed (purple dot next to the name).

The vent plan is as follows:

Open RV2

Open VM3

Open V7

Open V4

Shut down TP2

Install new gauges

Will add to post with updates post vent.

Attachment 1: Screenshot_2022-04-12_08-42-33.png
Screenshot_2022-04-12_08-42-33.png
Attachment 2: 81CB0936-1B19-4722-8A32-C3DC1D1FBC21.heic
Attachment 3: 2262ECEF-6200-4E95-8E32-C83CB9EB4F17.heic
Attachment 4: B5712B34-ECF6-43BE-BCB2-38CD775CF653.heic
Attachment 5: BEE14A07-976A-45C2-82A0-1774D377941E.heic
Attachment 6: 84557D6E-6AAE-47CF-A3AD-1DF329FEB550.heic
  16771   Mon Apr 11 21:44:14 2022 KojiSummaryBHDPart IIa of BHR upgrade - POY11 debugging

We took out the right most PD interface board D990543 and removed the 74LS04 chips. In fact two out of 4 were already replaced with enabling wires, however it seemed that one of the remaining two got failed.
These remaining two chips were removed and the enabling signals were connected to VCC (+5V). This operation made the status LED light not functional. (We didn't bother to fixed them by connecting them to the GND)

The new schematic diagram was attached here, and the corresponding DCC entry (https://dcc.ligo.org/D1900318-v1) was modified. Attachments #2-3 show the circuit after the changes, and the front view respectively.

Attachment 1: D990543A1.PDF
D990543A1.PDF D990543A1.PDF D990543A1.PDF
Attachment 2: PXL_20220412_023941730.jpg
PXL_20220412_023941730.jpg
Attachment 3: PXL_20220412_023950567.jpg
PXL_20220412_023950567.jpg
  16770   Mon Apr 11 21:13:21 2022 PacoSummaryBHDPart IIa of BHR upgrade - POY11 debugging

[Paco, Koji]

I asked Koji for some advice regarding closing the loop on YARM using POY11. A few things seemed off including

  • The YARM transmission; which was peaking at ~ 20 (typically, TRY is normalized so that under nominal input power conditions we see TRY in the range [0, 1])
  • The POY11 DCPD level was quite low; we expect a few tens of uW in this low power configuration where no more than 100 mW are going through IMC.

We looked at the POY11 RFPD first. We tried flashing an incandescent lamp in-situ and saw some weak response using an oscilloscope and the DC Out readout. We then used the OPHIR power meter and recorded ~ 1.2 uW of light incident on the POY11 RFPD... Initially, we suspected our beam was not filling the PD sensitive area (~ 2 mm diameter), but a quick estimate using the 200 mm focusing lens currently installed in the ITMY table gave us quite a generous margin of error... so we questioned the OPHIR measurement. We swapped the power meter and this time got ~ 18.4 uW, which is more in line with what we expected (phew).

Moving on, from the POY11 RFPD responsivity, and our ~ 20 uW of incident power, we expected on the order of a 20 mV of DC Output, but weren't really seeing this on the scope, so we decided to test the pins on the power of the RFPD. The DB15 cable not only supplies bipolar 15 VDC but also monitors several other test points such as Tsens, or DCout in the RFPD. We quickly noticed a weird signal on the ENAB testpoint, so we removed POY11 RFPD from the ITMY table and took it to the PD testbench. After redoing the soldering on the breakout board and RF amplifier (ZXX-500LN+), which we tested separately, we saw the expected behavior using a +- 15 VDC power supply... thus verifying that the RFPD and breakout board seemed to work ok. We turned our attention to the upstream DB15 connection, and after quickly checking the newly run cables, we ended up debugging the eurocrate PD interface. After attempting a simple power cycle and failing, we removed this card and looked at the schematic. It would seem that the logic enabling ICs (one or both) failed, thus preventing the card from enabling its outputs correctly.... We bypassed the logic by soldering pins 4,8,14 on U1, U20 and then checked the circuit in-situ, and we saw it worked fine again.

Now none of the status LEDs (which are driven by the logic IC portion) work on this card, but the card itself works fine at least for POY11.

We moved on, and installed the DB15 cables, checking the functionality at every step... Then we looked at the POY11_I_ERR signal and were happy to see nice pdh wavelets. We pushed forward a little bit more to try and lock YARM. First, we went to the Y end and centered the ETMY Oplev so as to register the position where the YARM is flashing... The ITMY Oplev is still not online. Then, we optimized the ETMY damping gains somewhat to try and make it less noisy, and finally, played with the LSC YARM loop gain to attempt locking. This last push was not as successful, but we have an idea of what next steps are needed to reduce the SUS noise, including

  • Install ITMY Oplevs, and close loop
  • Optimize ITMY damping gains

To be continued....

ELOG V3.1.3-